The percentage of microorganisms remaining vs. exposure time to autoclave sterilization process follows % = 70-t*2. The percentage of microorganisms
remaining vs. exposure time to dry heat sterilization process follows % = 80-t*5. At what time both the methods are equally efficient?
1
Expert's answer
2020-04-22T09:22:02-0400

For both the process to be equally efficient it must be needed that percentage for both should be equal.So, we can conclude that,

"70-t\times2=80-5\times t"

"3t=10"

t=3.33 sec.

So, for this time it will be equally efficient